Let a, b, q, r ∈ Z and suppose that a = bq + r. Show that gcd(a,b) = gcd(b,r)
1
Expert's answer
2012-08-28T10:09:33-0400
We use Bezout identity: d=gcd(a,b)=au+bv , for some u,v.
Then
d=(bq+r)u+bv=b(qu+v)+ru. Last equality shows that gcd(b,r)=d=gcd(a,b)